# D-LEAF: Localizing and Correcting Hallucinations in Multimodal LLMs via Layer–to-head Attention Diagnostics

### v1

@@ 4.2 Attention Head Localization
After pinpointing abnormal layers, we must identify the specific attention heads within them that require modification. Motivated by the “text inertia” phenomenon (Liu et al., 2024b), we introduce Image Attention Focus (IAF), a metric that quantifies the extent to which each attention head attends to image tokens (i.e., visual regions).
@@
Similarly, to validate the effectiveness of these two metrics, we repeated the Wilcoxon signed‐rank test described above in MiniGPT-4. As shown in Figure 4 right, we plotted the distributions and found that, at p << 0.001, every attention head exhibits significantly higher IAF for real tokens than for hallucinated ones.
Analogous to LIAE, to validate the soundness of our metric, we also evaluate Image Attention Entropy (IAE) and a hybrid of IAF and IAE for identifying and correcting anomalous attention heads; implementation details are provided in Appendix D.2. Across multiple model architectures, however, we find that using IAF alone consistently achieves the strongest hallucination suppression.
We also verified the reliability of our metrics on other model architectures, with specific results provided in Appendix D.3.
@@ 4.4 Dynamic Layer-wise Entropy and Attention Fusion
BAS denotes the score of the best-performing layer encountered so far in the forward pass, and it is updated incrementally. During inference, if the current layer’s LIAE is lower than BAS, we set BAS equal to LIAE; if LIAE exceeds BAS, we deem the attention heads in this layer to be underperforming and therefore in need of correction. We then rank all heads in the current layer by their IAF values, select the $n$ worst-performing heads, and update their attention matrices according to equation 6.
Our dynamic localization‑and‑correction algorithm not only suppresses hallucinations at a finer granularity within the model’s architectural layers but also leverages inter‑layer relationships to pinpoint anomalies with greater accuracy. The detailed workflow of the algorithm is shown in Figure 5, and its pseudo-code is provided in the Appendix D.5.
Figure 5: Architecture of our D-LEAF.
@@ 5.1 Experimental Settings
We conducted a detailed evaluation of our proposed algorithm using three established benchmarks: CHAIR Rohrbach et al. (2018), POPE Li et al. (2023b), and MMHal‑Bench Sun et al. (2023). These benchmarks were used to assess the effectiveness of our method in suppressing model hallucinations. Detailed descriptions of the benchmarks are provided in Appendix B.1.
@@
We evaluate our method D-LEAF on 5 models: LLaVA-1.5 (7B) (Liu et al., 2024a), MiniGPT-4 (Zhu et al., 2023), Shikra (Chen et al., 2023), InstructBLIP (Dai et al., 2023) and Qwen-VL Bai et al. (2023). And we compare it with several existing SOTA training-free hallucination suppression algorithms, including Greedy Search and Nuclear Sampling, SPIN (Sarkar et al., 2025), PAI (Liu et al., 2024b), VCD (Leng et al., 2024), and DAMRO (Gong et al., 2024). More details are provided in Appendix B.2.
@@
The lower the CI and CS values, the better the effect of hallucination suppression. However, suppressing hallucinations can affect the model’s outputs; therefore, we also employ Accuracy and F1 Score to rigorously evaluate the correctness of its generated answers.
@@
Greedy 47.08 $\pm$ 1.54 13.00 $\pm$ 0.59 34.00 $\pm$ 1.98 10.82 $\pm$ 0.59 54.64 $\pm$ 2.84 14.96 $\pm$ 1.37 48.12 $\pm$ 2.98 14.18 $\pm$ 1.12 46.88 $\pm$ 1.56 12.72 $\pm$ 0.36
Sampling 53.44 $\pm$ 2.21 16.30 $\pm$ 1.26 33.80 $\pm$ 1.17 11.78 $\pm$ 0.62 57.10 $\pm$ 1.80 16.14 $\pm$ 0.88 47.04 $\pm$ 1.47 13.24 $\pm$ 0.77 47.04 $\pm$ 1.47 13.24 $\pm$ 0.77
VCD 55.38 $\pm$ 1.17 15.20 $\pm$ 1.21 55.16 $\pm$ 2.25 14.96 $\pm$ 1.26 51.40 $\pm$ 2.21 13.62 $\pm$ 0.75
PAI 35.28 $\pm$ 1.69 9.46 $\pm$ 0.67 27.92 $\pm$ 1.47 10.06 $\pm$ 0.89 54.64 $\pm$ 2.51 14.02 $\pm$ 1.49 59.24 $\pm$ 2.00 16.10 $\pm$ 0.67 47.64 $\pm$ 2.26 12.92 $\pm$ 0.68
DAMRO 46.44 $\pm$ 1.66 12.78 $\pm$ 0.60
SPIN 29.04 $\pm$ 2.46 8.70 $\pm$ 0.55 24.56 $\pm$ 1.62 9.40 $\pm$ 1.74 38.56 $\pm$ 2.35 10.88 $\pm$ 0.56 48.80 $\pm$ 2.65 14.04 $\pm$ 1.23 33.72 $\pm$ 2.54 9.42 $\pm$ 0.48
D-LEAF 23.44 $\pm$ 2.63 6.72 $\pm$ 0.49 11.56 $\pm$ 1.69 4.72 $\pm$ 0.95 26.35 $\pm$ 1.32 10.62 $\pm$ 0.87 22.44 $\pm$ 2.75 8.48 $\pm$ 5.93 25.24 $\pm$ 1.55 7.96 $\pm$ 1.65
@@
Figure 6: MMHal-Bench Evaluation on LLaVA, MiniGPT-4 and Shikra.
@@ 5.2 Main Results
We use MMHal-Bench and GPT-4 assist to evaluate the performance of D-LEAF in more complex scenarios. From Figure 6, the experimental results indicate that our method could achieve better results in all three models, especially in LLaVA 7B. While for more image-based question types, like attributes and adversarial objects, our method did not achieve a noticeable improvement in MiniGPT-4 and Shikra. In the average performance across the eight evaluation dimensions, there is a certain degree of improvement compared to the baseline after incorporating D-LEAF.
@@
To evaluate whether our algorithm maintains real-time efficiency without incurring significant throughput loss, we measured the token-per-second generation rate of each model under different algorithms, as shown in Figure 7. Our method showed the least reduction in throughput compared to the baseline. It outperformed other state-of-the-art techniques, including attention-head correction methods such as SPIN and PAI.
In Appendix E, we provide visualizations across diverse MLLMs to further present instances of hallucination corrections by our method.
Figure 7: Throughput comparison with existing methods given by the number of tokens generated per second in LLaVA (left), MiniGPT-4 (middle) and Shikra (right).
@@ 5.3 Ablation Study
[…]
In addition, we conducted ablations on the mixing coefficient $\gamma$ and the number of suppressed attention heads $n$, as shown in Table 4. The results demonstrate the strong robustness of our method: across a wide range of $\gamma$ values (0.03 to 0.98), our algorithm consistently outperforms the baseline. We further observe that increasing $\gamma$ leads to a significant improvement in hallucination suppression across all three models. However, overly large $\gamma$ values result in a drop in F1 score, whereas appropriately chosen $\gamma$ achieves a favorable balance—reducing hallucinations while maintaining high F1 performance.
Building on the above, we further investigate the impact of the number of corrected attention heads, $n$, on hallucination reduction. We find that in models such as Instructblip, which leverage a learnable querying transformer to establish vision–language connections with only 32 image tokens as LVLM input, correcting even a small subset of attention heads achieves strong suppression performance. However, as the number of corrected heads increases, the model’s output capability deteriorates significantly: the F1 score drops sharply and the generated responses become markedly shorter. We attribute this to excessive correction disrupting the model’s normal output dynamics, causing it to prefer shorter responses as a way of avoiding hallucinated tokens. Interestingly, contrary to this trend, in LLaVA, which employs an MLP to map the vision branch output into 576 image embeddings, the number of corrected attention heads does not substantially affect model performance.
Model $L$ CS CI F1
LLaVA baseline 47.08 $\pm 1.54$ 13.00 $\pm 0.59$ 77.22 $\pm 0.66$
with L 23.44 $\pm 2.63$ 6.72 $\pm 0.49$ 74.88 $\pm 1.51$
without L 26.20 $\pm 2.63$ 8.30 $\pm 1.32$ 74.28 $\pm 0.81$
MiniGPT-4 baseline 34.00 $\pm 1.98$ 10.82 $\pm 0.59$ 69.60 $\pm 0.99$
with L 11.56 $\pm 1.69$ 4.72 $\pm 0.95$ 66.06 $\pm 0.99$
without L 14.60 $\pm 1.51$ 6.76 $\pm 1.13$ 66.52 $\pm 0.61$
Instrutblip baseline 48.12 $\pm 2.98$ 14.18 $\pm 1.12$ 73.92 $\pm 0.82$
with L 22.28 $\pm 1.32$ 7.82 $\pm 2.88$ 70.06 $\pm 0.80$
without L 22.44 $\pm 2.75$ 8.48 $\pm 5.93$ 70.74 $\pm 1.13$
Table 4: Ablation Study of $\gamma$ and $n$. The best result is highlighted in bold, and the second-best is underlined.
Model $\gamma$ CS CI F1
LLaVA - 47.8 13.3 76.2
0.03 27.2 10.0 72.1
0.30 26.2 10.1 71.9
0.50 31.6 9.9 76.1
0.90 24.6 7.5 75.6
0.98 20.6 6.2 75.3
MiniGPT-4 - 36.0 11.6 68.4
0.03 35.2 13.7 69.8
0.30 34.2 14.4 69.4
0.50 31.0 10.2 69.2
0.90 12.6 5.6 67.6
0.98 12.4 4.6 62.3
Instructblip - 53.4 15.1 72.7
0.03 50.6 15.6 73.7
0.30 48.6 14.8 73.5
0.50 46.0 13.2 74.2
0.90 14.8 4.8 67.1
0.95 10.6 9.1 65.3
Model $n$ CS CI F1
LLaVA - 47.8 13.3 76.2
5 28.8 8.2 76.9
10 28.6 7.1 76.4
15 24.8 7.4 75.4
20 25.2 5.0 75.0
25 20.6 6.2 75.3
MiniGPT-4 - 36.0 11.6 68.4
5 17.2 14.9 66.9
10 12.6 5.4 67.3
15 9.6 4.7 59.5
20 6.8 9.2 54.8
25 5.4 1.5 52.7
Instructblip - 53.4 15.1 72.7
5 23.0 4.4 70.5
10 1.4 20.2 26.5
15 1.2 3.5 39.2
20 2.8 1.6 33.2
25 1.8 8.1 25.3
@@ Appendix A Structure of The Appendix
Appendix B describes the datasets used in our validation experiments as well as the baseline methods for comparison.
@@
Appendix E provides qualitative case studies showcasing high-quality answers generated by models after applying our algorithm.
@@ B.2 Baselines
SPIN (Sarkar et al., 2025) and PAI (Liu et al., 2024b): The latest SOTA approach leverages attention‑head mechanisms to effectively suppress hallucinations.
@@
VCD (Leng et al., 2024): A technique that introduces noise into images to create amateur models for contrastive decoding.
@@
DAMRO (Gong et al., 2024): This method leverages the ViT’s CLS token to selectively filter out high‑attention background outliers and eliminate their influence during decoding.
@@ Appendix C Preliminary
In this section, we provide a detailed overview of the analysis tools employed, such as LogitLens and the Wilcoxon signed-rank test.
C.1 LogitLens
LogitLens (nostalgebraist, 2020) is an interpretability technique that directly maps each hidden state $x^{l}$ to the model’s vocabulary distribution by first applying the LayerNorm transformation and then projecting through the unembedding matrix $W_{U}$, as shown in equation 10.

### v2

@@ 4.2 Attention Head Localization
After pinpointing abnormal layers, we must identify the specific attention heads within them that require modification. Motivated by the “text inertia” phenomenon [liu2407paying], we introduce Image Attention Focus (IAF), a metric that quantifies the extent to which each attention head attends to image tokens (i.e., visual regions).
@@
Figure 6: IAF Distribution across object token types in the MiniGPT-4.
Similarly, to validate the effectiveness of these two metrics, we repeated the Wilcoxon signed‐rank test described above in MiniGPT-4. As shown in Fig. 6, we plotted the distributions and found that, at p << 0.001, every attention head exhibits significantly higher IAF for real tokens than for hallucinated ones.
Analogous to LIAE, to validate the soundness of our metric, we also evaluate Image Attention Entropy (IAE) and a hybrid of IAF and IAE for identifying and correcting anomalous attention heads; implementation details are provided in Sec. D.2. Across multiple model architectures, however, we find that using IAF alone consistently achieves the strongest hallucination suppression.
We also verified the reliability of our metrics on other model architectures, with specific results provided in Sec. D.3.
@@ 4.4 Dynamic Layer-wise Entropy and Attention Fusion
BAS denotes the best score of the layer encountered so far in the forward pass, and it is updated incrementally. During inference, if the current layer’s LIAE is lower than BAS, we set BAS equal to LIAE; if LIAE exceeds BAS, we deem the attention heads in this layer to be underperforming and therefore in need of correction. We then rank all heads in the current layer by their IAF values, select the $n$ worst-performing heads, and update their attention matrices according to Eq. 4.
Our dynamic localization‑and‑correction algorithm not only suppresses hallucinations at a finer granularity within the model’s architectural layers but also leverages inter‑layer relationships to pinpoint anomalies with greater accuracy. The detailed workflow of the algorithm and its pseudo-code is provided in the Sec. D.5.
5 Theoretical Analysis of D-LEAF
Inspired by previous studies [uppaal2024model, yang2025nullu], in this section we provide a theorem to analyze the connection between D-LEAF and DPO from the perspective of optimization objectives. This analysis establishes the connection between D-LEAF and DPO, and a more detailed discussion is provided in Appendix E.
Our subsequent analysis is derived by examining a simple logistic model $\pi_{\text{W}}$ for the output conditioned on the input $x$. The conditional probability can be expressed as:
$\pi_{\text{W}}(y|x_{i})=\text{Z}_{\text{W}}^{-1}\text{exp}(o_{y}^{\top}\text{W}x_{i})$ (6)
where $w_{y}$ is the output decoding vector for any $y\in\text{Vocabulary}$, $\text{Z}_{\text{W}}$ is the normalization factor.
Theorem 1 (Equivalence of Optimization Directions in DPO and D-LEAF).
Let $x_{i}^{+}$ and $x_{i}^{-}$ denote the hallucinated and preferred embeddings at any layer of an MLLM, respectively. The gradient of the DPO objective satisfies
$\nabla_{\!W}L_{\mathrm{DPO}}\propto\sum_{i=1}^{N}\left(o_{y_{i}^{+}}(x_{i}^{+}-x_{i}^{-})^{\top}+(o_{y_{i}^{+}}-o_{y_{i}^{-}})(x_{i}^{-})^{\top}\right).$
Thus, DPO reduces the feature difference $x_{i}^{+}-x_{i}^{-}$ by pushing the hallucinated embedding toward the preferred one.
In D-LEAF, attention-head fusion modifies the attention matrix so that the updated embedding
$\tilde{x}^{l}=\mathrm{FFN}(\mathrm{MHA}_{\gamma}(x^{l-1}))$
decreases the same feature discrepancy $\|\tilde{x}^{l}-x_{i}^{-}\|$ through attention reweighting.
Therefore, both DPO and D-LEAF optimize toward the same descent direction by reducing the embedding gap between the hallucinated and preferred outputs.
Remark 1.
Theorem 1 reveals that D-LEAF is not merely a heuristic attention-correction mechanism. Instead, its update rule can be interpreted as an implicit reparameterization of the DPO gradient in embedding space. Specifically, while DPO explicitly enforces preference alignment by reducing the feature discrepancy $x_{i}^{+}-x_{i}^{-}$ through gradient descent, D-LEAF achieves an equivalent effect by modulating the attention matrices that determine each layer’s embedding. Consequently, D-LEAF performs a training-free approximation of the DPO optimization objective, which explains why D-LEAF effectively suppresses hallucinations despite incurring no additional fine-tuning cost.
@@ 6.1 Experimental Settings
We conducted a detailed evaluation of our proposed algorithm using three established benchmarks: CHAIR [rohrbach2018object], POPE [li2023evaluating], and MMHal‑Bench [sun2023aligning]. These benchmarks were used to assess the effectiveness of our method in suppressing model hallucinations. Detailed descriptions of the benchmarks are provided in Sec. B.1.
@@
We evaluate our method D-LEAF on 5 models: LLaVA-1.5 (7B) [liu2024improved], MiniGPT-4 [zhu2023minigpt], Shikra [chen2023shikra], InstructBLIP [dai2023instructblip] and Qwen-VL [Qwen-VL]. And we compare it with several existing SOTA training-free hallucination suppression algorithms, including Greedy Search and Nuclear Sampling, SPIN [sarkar2025mitigating], PAI [liu2407paying], VCD [leng2024mitigating], and DAMRO [gong2024damro]. More details are provided in Sec. B.2.
@@
Greedy 47.08 ±1.54 13.00 ±0.59 34.00 ±1.98 10.82 ±0.59 54.64 ±2.84 14.96 ±1.37 48.12 ±2.98 14.18 ±1.12 46.88 ±1.56 12.72 ±0.36
Sampling 53.44 ±2.21 16.30 ±1.26 33.80 ±1.17 11.78 ±0.62 57.10 ±1.80 16.14 ±0.88 47.04 ±1.47 13.24 ±0.77 47.04 ±1.47 13.24 ±0.77
VCD 55.38 ±1.17 15.20 ±1.21 55.16 ±2.25 14.96 ±1.26 51.40 ±2.21 13.62 ±0.75
PAI 35.28 ±1.69 9.46 ±0.67 27.92 ±1.47 10.06 ±0.89 54.64 ±2.51 14.02 ±1.49 59.24 ±2.00 16.10 ±0.67 47.64 ±2.26 12.92 ±0.68
DAMRO 46.44 ±1.66 12.78 ±0.60
SPIN 29.04 ±2.46 8.70 ±0.55 24.56 ±1.62 9.40 ±1.74 38.56 ±2.35 10.88 ±0.56 48.80 ±2.65 14.04 ±1.23 33.72 ±2.54 9.42 ±0.48
D-LEAF 23.44 ±2.63 6.72 ±0.49 11.56 ±1.69 4.72 ±0.95 26.35 ±1.32 10.62 ±0.87 22.44 ±2.75 8.48 ±5.93 25.24 ±1.55 7.96 ±1.65
@@
Figure 7: MMHal-Bench Evaluation on LLaVA.
@@ 6.2 Main Results
We use MMHal-Bench and GPT-4 assist to evaluate the performance of D-LEAF in more complex scenarios. From Fig. 7, the experimental results indicate that our method could achieve better results in in LLaVA 7B. However, for comparison-type tasks involving multi-object attribute reasoning, the model’s performance is relatively weak. We also provide the results of MiniGPT-4 and Shikra on MMHal-Bench in Sec. B.3, where our method consistently improves upon the baseline performance on average.
Figure 8: Throughput comparison with existing methods in LLaVA.
@@
To evaluate whether our algorithm maintains real-time efficiency without incurring significant throughput loss, we measured the token-per-second generation rate of LLaVA under different algorithms, as shown in Fig. 8. Our method showed the least reduction in throughput compared to the baseline. It outperformed other state-of-the-art techniques, including attention-head correction methods such as SPIN and PAI. We repeated this experiment on MiniGPT-4 and Shikra and observed consistent results, as detailed in Sec. B.4.
In Appendix F, we provide visualizations across diverse MLLMs to further present instances of hallucination corrections by our method.
@@ 6.3 Ablation Study
D-LEAF incorporates two primary hyperparameters: $\gamma$ and the number of heads $n$. We present the effect of varying two parameters on CHAIRS in the main text, as illustrated in Fig. 9. The results demonstrate the strong robustness of our method: across a wide range of $\gamma$ (0.03 to 0.98) and $n$ values (5 to 25), our algorithm consistently outperforms the baseline. In the Fig. 12, we further provide the impact of hyperparameter variations on CHAIRI and F1, along with a more detailed analysis of the results. Moreover, to ensure the completeness of our study, we also examine whether restricting the D-LEAF algorithm to specific layers yields additional gains. The results in Fig. 12 and Sec. D.4 demonstrate that, unlike PAI [liu2407paying] and Deco [wang2024mllm], D-LEAF consistently achieves significant suppression of hallucinations regardless of whether the layer prior is applied.
Figure 9: Ablation Study results for hyperparamter $\gamma$ and $n$.
@@ Appendix A Structure of The Appendix
Appendix B details the datasets used in our validation experiments as well as the baseline methods for comparison. It also provides additional experiments that complement the main text, including throughput estimates across different models, ablation studies, and related analyses.
@@
Appendix F provides qualitative case studies showcasing high-quality answers generated by models after applying our algorithm.
Appendix E provides the derivation of the DPO optimization objective and further elaborates on the conceptual connection between D-LEAF and DPO.
@@ B.2 Baselines
SPIN [sarkar2025mitigating] and PAI [liu2407paying]: The latest SOTA approach leverages attention‑head mechanisms to effectively suppress hallucinations.
@@
VCD [leng2024mitigating]: A technique that introduces noise into images to create amateur models for contrastive decoding.
@@
DAMRO [gong2024damro]: This method leverages the ViT’s CLS token to selectively filter out high‑attention background outliers and eliminate their influence during decoding.
@@
B.3 MMHal-Bench Evaluation
We present in Fig. 10 the performance of D-LEAF compared with the baseline on MiniGPT-4 and Shikra for the MMHal-Bench tasks. D-LEAF shows substantial improvement in descriptive tasks such as the holistic description task, and it also achieves a noticeable gain in the average performance across all task categories.
Figure 10: MMHal-Bench Evaluation on MiniGPT-4 and Shikra.
B.4 Throughput Estimation
In the main text, we demonstrated that on LLaVA our algorithm achieves throughput closest to greedy search among all methods. To further validate its effectiveness, we measured throughput on MiniGPT-4 and Shikra. As shown in Fig. 11, our approach still incurs the smallest throughput degradation while maintaining high hallucination suppression rates and preserving output detail.
Figure 11: Throughput comparison with existing methods given by the number of tokens generated per second in MiniGPT-4 (left) and Shikra (right).
B.5 Ablation Study
[…]
Model $L$ CS CI F1
LLAVA baseline 47.08 $\pm 1.54$ 13.00 $\pm 0.59$ 77.22 $\pm 0.66$
with L 23.44 $\pm 2.63$ 6.72 $\pm 0.49$ 74.88 $\pm 1.51$
without L 26.20 $\pm 2.63$ 8.30 $\pm 1.32$ 74.28 $\pm 0.81$
MiniGPT-4 baseline 34.00 $\pm 1.98$ 10.82 $\pm 0.59$ 69.60 $\pm 0.99$
with L 11.56 $\pm 1.69$ 4.72 $\pm 0.95$ 66.06 $\pm 0.99$
without L 14.60 $\pm 1.51$ 6.76 $\pm 1.13$ 66.52 $\pm 0.61$
Instrutblip baseline 48.12 $\pm 2.98$ 14.18 $\pm 1.12$ 73.92 $\pm 0.82$
with L 22.28 $\pm 1.32$ 7.82 $\pm 2.88$ 70.06 $\pm 0.80$
without L 22.44 $\pm 2.75$ 8.48 $\pm 5.93$ 70.74 $\pm 1.13$
In addition to the analysis in the main text on the impact of the mixing coefficient $\gamma$ and the number of suppressed attention heads $n$ on $\mathrm{CHAIR}_{S}$, we further report here the influence of these hyperparameters on $\mathrm{CHAIR}_{I}$ and F1, as shown in Fig. 12 and Fig. 13. The results demonstrate the strong robustness of our method: across a wide range of $\gamma$ values (0.03 to 0.98), our algorithm consistently outperforms the baseline. We further observe that increasing $\gamma$ leads to a significant improvement in hallucination suppression across all three models. However, overly large $\gamma$ values result in a drop in F1 score, whereas appropriately chosen $\gamma$ achieves a favorable balance—reducing hallucinations while maintaining high F1 performance.
As for the impact of the number of corrected attention heads, $n$, on hallucination reduction, we find that in models such as InstructBLIP, which leverage a learnable querying transformer to establish vision–language connections with only 32 image tokens as MLLM input, correcting even a small subset of attention heads achieves strong suppression performance. However, as the number of corrected heads increases, the model’s output capability deteriorates significantly: the F1 score drops sharply and the generated responses become markedly shorter. We attribute this to excessive correction disrupting the model’s normal output dynamics, causing it to prefer shorter responses as a way of avoiding hallucinated tokens. Interestingly, contrary to this trend, in LLaVA, which employs an MLP to map the vision branch output into 576 image embeddings, the number of corrected attention heads does not substantially affect model performance.
Figure 12: Ablation Study results with CHAIRI for hyperparamter $\gamma$ and $n$.
Figure 13: Ablation Study results with F1 for hyperparamter $\gamma$ and $n$.
@@ Appendix C Preliminary
In this section, we provide a detailed overview of the analysis tools employed, such as LogitLens and the Wilcoxon signed-rank test, and the forward process of MLLMs.
C.1 Model Forward Process
The language decoder module here consists of multiple transformer components [vaswani2017attention]. Each transformer block comprises two sublayers: a multi-head attention (MHA) mechanism and a feed-forward network (FFN). MHA begins by taking the combined text and image embedding vectors $X\in R^{N\times d}$ as input, projecting them into the query $(Q)$, key $(K)$, and value $(V)$ spaces, and then computing the output of the MHA module. The output of MHA is fed into the FFN module, and the final output of the current encoder block is obtained through residual flow, as shown in Eq. 9 and Eq. 10.
$x_{n}^{mid,l}=\sum_{h=1}^{H}\text{Attn}^{(l,h)}(X_{\leq n}^{l-1})W_{U}$ (9)
C.2 LogitLens
LogitLens [logitlens] is an interpretability technique that directly maps each hidden state $x^{l}$ to the model’s vocabulary distribution by first applying the LayerNorm transformation and then projecting through the unembedding matrix $W_{U}$, as shown in Eq. 11.